# Sample Compression for Continual Learning

### v1

@@ 3.2 Compression and Reconstruction Scheme
In line with the prevalent literature, we presented the sample compression framework (Section 2.2) based on a compression function providing a compression set and optionally a message. In the case of CoP2L, the reconstruction scheme relies on two compression sets $(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}})$ and a message pair $(\mu_{1},\mu_{2})$, which we explain later in this section. [^1] The idea of using multiple compression sets appeared in Marchand & Shawe-Taylor (2002) and Marchand et al. (2003), but in a setting without a message Two challenges prevent CoP2L from being used as its own compression and reconstruction function as the original P2L.
The first challenge comes from the sampling of the buffer ${\mathcal{B}}$. When training using CoP2L, datapoints from the buffer can be chosen to be part of the compression set. At the next task, these datapoints may still be available to train the model (if they are not excluded from the buffer by the sampling step). Therefore, the reconstruction function needs to know when to remove the datapoint. The message thus provide the task after which the datapoint was removed during CoP2L execution. As the reconstruction function only requires a message for datapoints who were removed from the buffer after being added to the compression set $S^{\operatorname{\mathbf{i}}}$, a second compression set $S^{\operatorname{\mathbf{j}}}$ is dedicated to datapoints that necessitate a message. The message $\mu_{1}$ given is chosen among $\{2,\ldots,T\}$ for each datapoint belonging to $S^{\operatorname{\mathbf{j}}}$.
The second one is the use of the bound as stopping criterion in mP2L. Recall that the reconstruction function does not have access to the whole dataset. Thus, it cannot recover the bound value computed during the initial training phase (see Line 11 of Algorithm 2) and use it as a stopping criterion. To address this inconvenience, the number of iterations to perform is provided to mP2L as a message $\mu_{2}=(\mu_{2}^{1},\ldots,\mu_{2}^{T})$. That is, for each task $t$, with a buffer $\mathcal{B}$, the message component $\mu_{2}^{t}$ is chosen in $\{1,\ldots,n_{t}+|\mathcal{B}|\}$, giving the number of mP2L’s iterations to perform.
$\mathscr{M}(\operatorname{\mathbf{j}},T)\,=\,\{2,\ldots,T\}^{|\operatorname{\mathbf{j}}|}\times\left[\bigtimes\limits_{t=1}^{T}\{1,\ldots,n_{t}+|{\mathcal{B}}|\}\right].$ (9)
Reconstruction function.
Under the assumption that the input parameters $\theta_{0},\gamma,m,\omega$ are the same for the compression function and for CoP2L, the reconstruction function must output the exact same predictor as CoP2L. However, it only has access to the compression sets and the messages. The first challenge of the compression function is addressed using $\mu_{1}$ from line 6 to 11 of Algorithm 5. For each datasets, we verify if a datapoint was previously excluded from the buffer by the sampling function. If so, we remove it from the buffer. The second challenge is adressed in line 4 of Algorithm 5 by giving the message $\mu_{2}$ to mP2L as stopping criterion.
Algorithm 4 Compression function of CoP2L
@@
0:  $S_{1},S_{2},\ldots,S_{T}$Training sets}
1:  $B_{i}\leftarrow\emptyset,C_{i}\leftarrow\emptyset,C^{\mu}_{i}\leftarrow\emptyset,\mu_{1,i}\leftarrow\emptyset\quad\forall i=1,\ldots,T$
@@
7:     $C_{i}\leftarrow C_{i}\cup(C^{\star}\cap S_{i})\quad\forall i=1,\ldots,t$
@@
12:           $C^{\mu}_{i}\leftarrow C^{\mu}_{i}\cup\{({\mathbf{x}}_{i,j},y_{i,j})\}$
13:           $C_{i}\leftarrow C_{i}\setminus\{({\mathbf{x}}_{i,j},y_{i,j})\}$
@@
19:  return $\theta_{T},\{C_{t}\}_{t=1}^{T},\{C^{\mu}_{t}\}_{t=1}^{T},\{M_{t}\}_{t=1}^{T},N$. {Learned parameters, compression sets and messages sets}
Algorithm 5 Reconstruction function of CoP2L
@@
0:  $C_{1},C_{2},\ldots,C_{T},\ C^{\mu}_{1},C^{\mu}_{2},\ldots,C^{\mu}_{T}$Compression sets}
0:  $\mu_{1,1},\mu_{1,2},\ldots,\mu_{1,T},\,\mu_{2}${Message sets}
@@
3:     $\hat{S_{t}}\leftarrow\{({\mathbf{x}},y,1)\}_{({\mathbf{x}},y)\in C_{t}}$
@@
5:     $B_{t}\leftarrow(C_{t}\setminus C^{\star})\cup C^{\mu}_{t}$
@@
7:        for $(\boldsymbol{x}_{i,j},y_{i,j},\omega_{i})\in C^{\mu}_{i}$ do
@@
3.3 Generalization Bound
The last subsection provided the compression and reconstruction functions of CoP2L. Thus, we know CoP2L provides a sample-compressed predictor. Theorem 2.1 can be used to obtain generalization bounds on the last task, which we do in mP2L when we compute the bound $\Psi$, but it holds for only one distribution of data. Therefore, Theorem 2.1 cannot be applied to bound all tasks learned by our continual learning scheme.
Le us now consider a reconstruction function that takes as input two compression sets $\operatorname{\mathbf{i}},\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n)$ and a message $\mu=(\mu_{1},\mu_{2})\in\mathscr{M}(\operatorname{\mathbf{j}},T)$. For any distribution $\operatorname{\mathcal{D}}_{t}$ and any dataset $S_{t}\sim\operatorname{\mathcal{D}}_{t}$, we denote the loss on both compression sets as ^$\widehat{{\mathcal{L}}}_{S_{t}}^{\ \operatorname{\mathbf{i}}}(\theta)$ and ^$\widehat{{\mathcal{L}}}_{S_{t}}^{\ \operatorname{\mathbf{j}}}(\theta)$. The loss on the joint compression sets is denoted ^$\widehat{{\mathcal{L}}}_{S_{t}}^{\ \operatorname{\mathbf{i}}\cup\operatorname{\mathbf{j}}}(\theta)$ and the loss on the complement set of the joint set $\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}$ is denoted ^$\widehat{{\mathcal{L}}}_{S_{t}}^{\ \operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\!(\theta)$. Moreover, we denote the reconstruction function of CoP2L as $\operatorname{\mathscr{R}}_{1:T}(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu)=\operatorname{\mathscr{R}}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu;S_{1},\ldots,S_{t-1},S_{t+1},\ldots,S_{T}).$ This formulation is important for the following theorem, as the reconstruction function should be fixed for all datasets $S_{1}$ to $S_{T}$, with the exception of $S_{t}$.
Theorem 3.1.
$\forall t\in\{1,\ldots,T\},{\mathbf{i}},\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n_{t}),\mu\in\mathscr{M}(\operatorname{\mathbf{j}},T):$
$\leq\frac{1}{n_{t}-|{\mathbf{i}}|-|\mathbf{j}|}\bigg{[}\log\matrixquantity(n_{t}\\ |{\mathbf{i}}|)+\log\matrixquantity(n_{t}-\operatorname{|\mathbf{i}|}\\ |\mathbf{j}|)$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}+\sum_{i=1}^{T}\log\frac{1}{\zeta(\mu_{2}^{i})}+|\mathbf{j}|\log(T-1)+\log\frac{T}{\zeta(|{\mathbf{i}}|)\zeta(|\operatorname{\mathbf{j}}|)\delta}\bigg{]}$
with $\mu=(\mu_{1},\mu_{2})$.
The proof is given in Appendix B. It first relies on a tighter version of Theorem 2.1, where the $2\sqrt{n-\operatorname{|\mathbf{i}|}}$ term is removed and a second compression set is added. Then, this new result is adapted to the continual learning setting.
For all tasks $1$ through $T$, this new theorem encodes a tradeoff similar to Theorem 2.1. A more accurate predictor will achieve a tighter bound. However, a simpler predictor will achieve a tighter bound.
In the sample compression theory, we define the complexity of the model using the size of the compression sets $\operatorname{\mathbf{i}}$ and $\operatorname{\mathbf{j}}$ and the probability of choosing a message. The probability of the messages $\mu_{1}$ and $\mu_{2}$ are functions of the size of $\operatorname{\mathbf{j}}$ and the number of tasks. Indeed, the probability of $\mu_{1}$ is independent of its content but decays when $|\operatorname{\mathbf{j}}|$ and $T$ grow larger. Moreover, the probability of $\mu_{2}$ depends on the size of the compression set outputted by mP2L at each task $T$. Thus, the larger the compression sets are, the smaller the probability of both messages will be and the complexity of the model grows when the compression sets grow.
In conclusion, the model that will achieve the best possible bound is a model that minimizes the loss on the dataset whilst still keeping $\operatorname{\mathbf{i}}$ and $\operatorname{\mathbf{j}}$ the smallest possible.
4 Experiments
Figure 1: Accuracy (above) and forgetting (below) over replay buffer size for (left column) MNIST dataset, (middle column) Fashion-MNIST dataset, (right column) EMNIST dataset. The orange triangles show the average performance for the corresponding box plot.
Figure 2: CoP2L bounds with respect to tasks on the MNIST and Fashion-MNIST datasets
Datasets We carry-out experiments on MNIST (LeCun et al., 2010), Fashion-MNIST (Xiao et al., 2017), and EMNIST (Cohen et al., 2017) datasets on class-incremental learning. We divide the tasks into increments of 2 classes. This yields 5 tasks for MNIST and Fashion-MNIST datasets, and 13 tasks for the EMNIST dataset.
Implementation Details We have used the Avalanche continual learning framework for our experimental setup (Carta et al., 2023). Avalanche is a popular toolkit for continual learning. We have used their class-incremental MNIST and Fashion-MNIST dataset implementations, as well as the replay baseline in order to ensure compatibility with the results in the literature. We implement a block version of mP2L, which adds $k$ examples to the compression set at a time, as described in Algorithm 2 of Paccagnan et al. (2024). For the experiments, we use $k=8$. We provide the code for our experimental results in an anonymous repository [^2] https://anonymous.4open.science/r/sample_compression_continual_learning-00C7..
Model architecture and training details We use a simple multilayer perceptron (MLP) architecture for all experiments (SimpleMLP class from the Avalanche toolkit), and train using the SGD optimizer, with a learning rate of 0.001 for CoP2L and 0.01 for Replay. For all experiments with the Replay method, we train the model 20 epochs on each tasks. In contrast, CoP2L follows a different training strategy, continuing until it reaches zero errors on the complement set. We add a weight of $\omega=15$ in the cross-entropy loss when training on previous tasks.
@@ 4 Experiments
Computation of the bound To compute the bound, we need to keep track of the size of the compression sets and of the messages. We thus implement the compression function as described in Algorithm 4. However, we need to adjust the code to account for mP2L adding $k$ datapoints to the compression set at each iterations. When computing the message $\mu_{2}$, at iteration $t$, $\mu_{2}^{t}$ is the number of iterations of the mP2L algorithm. Knowing that mP2L outputs a compression set $C^{\star}_{t}$, and adds $k$ datapoints to the compression set at each iteration, the number of iterations is $\mu_{2}^{t}=\frac{|C^{\star}_{t}|}{k}$. Finally, when computing the bound, we use the code provided by Viallard et al. (2021) to invert the $\mathrm{kl}$ divergence.
4.1 Metric Comparison over Buffer Size
In Figure 1, we compare the accuracy and forgetting obtained with different replay buffer sizes of (1000, 2000, 3000, 4000, 5000) samples for CoP2L and replay. This represents roughly 8% to 40% of the per-task dataset size, for MNIST and Fashion-MNIST, and 10% to 52% for EMNIST. For each replay buffer size, we report the variation of accuracy obtained at the end of training averaged over different seeds, and test accuracy for different tasks. Note that we denote the accuracy obtained on a task $t$, after training the model for $T$ tasks with $\text{Accuracy}_{T,t}$. The accuracy we report is defined as
$\text{Average Accuracy}_{T}=\frac{1}{s\cdot T}\sum_{i=1}^{s}\sum_{t=1}^{T}\text{Accuracy}_{T,t}^{i},$ (10)
where $i$ index is for different seeds, and $t$ is for the task index. In our experiments, we used $s=5$ seeds to measure the variability with respect to weight initialization. We also report the average forgetting at task $T$ which is defined as follows,
$\text{Average Forgetting}_{T}=$ (11)
$\frac{1}{(T-1)\cdot s}\sum_{t=1}^{T-1}\sum_{i=1}^{s}$ $\text{Accuracy}_{t,t}^{i}-\text{Accuracy}_{T,t}^{i},$
and can be calculated when $T>1$. Note that in Figure 1, we report the average accuracy obtained after the final task, such that $T=5$ for MNIST and Fashion-MNIST and $T=13$ for EMNIST.
4.2 Empirical Study of the CoP2L generalization bound
In Figure 2, we showcase the generalization bound of CoP2L. We observe that the bound values for the generalization error always remain larger than the actual error incurred after each task, and we observe that the reported bound values are non-trivial. We also observe that the bound values generally follow similar trends as the actual error we report on the test set, while the estimate for the bound is calculated exclusively on the training set. In Figure 2, we provide the bounds for the models trained with a buffersize of 1000, but we give the results for all buffersizes, and all three datasets in Appendix C.
4.3 Ablation on CoP2L loss terms
We have conducted an ablation study to understand the effect of additional weighting that counteracts class imbalance that arises in the continual learning setup when new tasks are introduced (Standard replay counteracts this by extending the batch size to have equal class representation). In the modified P2L algorithm (2), we incorporate an additional weighting term over classes to adapt P2L for class-balanced continual learning. In order to assess whether this idea is indeed helpful, we have run an experiment on the MNIST dataset with and without the additional weighting idea. In Figure 3, we show this experiment. We observe that for both accuracy (left panel) and forgetting (right panel), the inclusion of the additional loss weighting term significantly improves the results.
4.4 Sensitivity analysis with respect to the task ordering
In order to also assess the sensitivity of CoP2L with respect to ordering of the tasks, on MNIST and Fashion-MNIST experiments we have conducted an experiment where we randomly shuffle the task order. For both dataset, we have tested our algorithm with 5 different random task ordering.
For each task, we report the cumulative accuracy
[…]
where $\text{Accuracy}_{T,t}^{o}$ denotes the accuracy obtained at task $t$ after finishing training until task $T$, and the $o$ index denotes the task order. We have used $O=5$ different task orderings. We present the results of this experiment in Figure 4. We observe that on both MNIST and Fashion-MNIST datasets CoP2L is able outperform with more tasks. We also observe that the variability of the results decrase more significantly for CoP2L compared to replay.
5 Conclusions
In this paper, we have proposed CoP2L, an algorithm rooted in the sample compression theory. To the best of our knowledge, this is the first attempt in the literature to employ sample compression theory within the continual learning context. We provided sample compression bounds for CoP2L and verified empirically that they were non-vacuous. We furthermore showed that on three different datasets, our approach is able to outperform experience replay, which is known to be a strong continual learning baseline.
Acknowledgements
We also wish to thank Benjamin Leblanc for his help proof-reading the manuscript. Mathieu Bazinet is supported by a FRQNT B2X scholarship (343192). Pascal Germain is supported by the NSERC Discovery grant RGPIN-2020-07223. Cem Subakan is supported by NSERC Discovery grant RGPIN-2023-05759. This research was also enabled in part by support provided by the Digital Research Alliance of Canada.
Appendix A Tighter sample compression $\mathrm{kl}$ bound
@@
In this section, we prove a tighter $\mathrm{kl}$ bound that is not derived from Theorem 2.1 but proved in a very similar way. To do so, we start by providing the Chernoff test-set bound (Langford, 2005) for losses in $[0,1]$ (Foong et al., 2022).
Theorem A.1 ((Foong et al., 2022)).
@@
$p\leq\mathrm{kl}^{-1}\quantity(\frac{1}{n}\sum_{i=1}^{n}X_{i},\frac{1}{n}\log\frac{1}{\delta}).$
@@
We prove the complement of the expression in Theorem A.2. We choose $\delta_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}=P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{i}})P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})P_{\mathscr{M}(\operatorname{\mathbf{j}})}(\mu)\delta$.
$\operatorname*{\mathbb{P}}_{S\sim\operatorname{\mathcal{D}}^{n}}\quantity(\exists\operatorname{\mathbf{i}},\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n),\mu\in\mathscr{M}(\operatorname{\mathbf{j}}):\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu))>\mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S}^{\,\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\quantity(\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu)),\frac{1}{n-\operatorname{|\mathbf{i}|}-|\operatorname{\mathbf{j}}|}\log\frac{1}{\delta_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}}))$
$\leq\sum_{\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)}\sum_{\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n)}\operatorname*{\mathbb{P}}_{S\sim\operatorname{\mathcal{D}}^{n}}\quantity(\exists\mu\in\mathscr{M}(\operatorname{\mathbf{j}}):\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu))>\mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S}^{\,\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\quantity(\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu)),\frac{1}{n-\operatorname{|\mathbf{i}|}-|\operatorname{\mathbf{j}}|}\log\frac{1}{\delta_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}}))$ (13)
@@
$\leq\sum_{\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)}\sum_{\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n)}\sum_{\mu\in\mathscr{M}(\operatorname{\mathbf{j}})}\operatorname*{\mathbb{P}}_{S\sim\operatorname{\mathcal{D}}^{n}}\quantity(\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu))>\mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S}^{\,\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\quantity(\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu)),\frac{1}{n-\operatorname{|\mathbf{i}|}-|\operatorname{\mathbf{j}}|}\log\frac{1}{\delta_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}}))$ (14)
$\leq\sum_{\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)}\sum_{\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n)}\sum_{\mu\in\mathscr{M}(\operatorname{\mathbf{j}})}\operatorname*{\mathbb{E}}_{S^{\operatorname{\mathbf{i}}}\sim\operatorname{\mathcal{D}}^{\operatorname{|\mathbf{i}|}}}\operatorname*{\mathbb{E}}_{S^{\operatorname{\mathbf{j}}}\sim\operatorname{\mathcal{D}}^{|\operatorname{\mathbf{j}}|}}\operatorname*{\mathbb{P}}_{S^{\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\sim\operatorname{\mathcal{D}}^{n-\operatorname{|\mathbf{i}|}-|\operatorname{\mathbf{j}}|}}\quantity(\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu))>\mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S}^{\,\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\quantity(\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu)),\frac{1}{n-\operatorname{|\mathbf{i}|}-|\operatorname{\mathbf{j}}|}\log\frac{1}{\delta_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}}))$ (15)
@@
$\leq\sum_{\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)}\sum_{\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n)}\sum_{\mu\in\mathscr{M}(\operatorname{\mathbf{j}})}\operatorname*{\mathbb{E}}_{S^{\operatorname{\mathbf{i}}}\sim\operatorname{\mathcal{D}}^{\operatorname{|\mathbf{i}|}}}\operatorname*{\mathbb{E}}_{S^{\operatorname{\mathbf{j}}}\sim\operatorname{\mathcal{D}}^{|\operatorname{\mathbf{j}}|}}\delta_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}$ (16)

### v3

@@ Appendix A Compression and Reconstruction Functions
In line with the prevalent literature, we introduced the sample compression framework (Section 2) based on a compression function providing a compression set and optionally a message. In the case of CoP2L, the reconstruction scheme relies on two compression sets $(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}})$ and a message pair $(\mu_{1},\mu_{2})$, which we explain later in this section. [^2] The idea of using multiple compression sets appeared in [40] and [39], but in a setting without a message.
Algorithm 4 Compression function of Continual Pick-To-Learn (CoP2L)
0:  $S_{1},S_{2},\ldots,S_{T}$Training sets}
@@
1:  $B_{t}\leftarrow\emptyset,S^{\operatorname{\mathbf{i}}}_{t}\leftarrow\emptyset,S^{\operatorname{\mathbf{j}}}_{t}\leftarrow\emptyset,\mu_{1,t}\leftarrow\emptyset\quad\forall t=1,\ldots,T$
@@
7:    $S^{\operatorname{\mathbf{i}}}_{i}\leftarrow S^{\operatorname{\mathbf{i}}}_{i}\cup(C^{\star}\cap S_{i})\quad\forall i=1,\ldots,t$
@@
12:          $S^{\operatorname{\mathbf{j}}}_{i}\leftarrow S^{\operatorname{\mathbf{j}}}_{i}\cup\{({\mathbf{x}}_{i,j},y_{i,j})\}$
13:          $S^{\operatorname{\mathbf{i}}}_{i}\leftarrow S^{\operatorname{\mathbf{i}}}_{i}\setminus\{({\mathbf{x}}_{i,j},y_{i,j})\}$
@@
19:  return $\theta_{T},\{S^{\operatorname{\mathbf{i}}}_{t}\}_{t=1}^{T},\{S^{\operatorname{\mathbf{j}}}_{t}\}_{t=1}^{T},\{\mu_{1,t}\}_{t=1}^{T},\mu_{2}$. {Learned parameters, compression sets and messages sets}
A.1 Compression Function
Two challenges prevent CoP2L (as presented by Algorithm 3) from being used as its own compression and reconstruction function. In this subsection, we expose each of these challenges while explaining how we the proposed compression algorithm (Algorithm 4) allows to circumvent them.
The first challenge comes from the sampling of the buffer $B^{\star}$ (Lines 6 to 8 of Algorithm 3). When training using CoP2L, datapoints from the buffer can be chosen to be part of the compression set. At the next task, these datapoints may still be available to train the model (if they are not excluded from the buffer by the sampling step). Therefore, the reconstruction function needs to know when to remove the datapoint. The message $\mu_{1}$ returned by Algorithm 4 thus provides the task index after which the datapoint was removed during CoP2L execution. As the reconstruction function only requires a message for datapoints who were removed from the buffer after being added to the compression set $S^{\operatorname{\mathbf{i}}}$, a second compression set $S^{\operatorname{\mathbf{j}}}$ is dedicated to datapoints that necessitate a message. The message $\mu_{1}$ given is chosen among $\{2,\ldots,T\}$ for each datapoint belonging to $S^{\operatorname{\mathbf{j}}}$.
The second challenge is the use of the bound as stopping criterion in mP2L. Recall that the reconstruction function does not have access to the whole dataset. Thus, it cannot recover the bound value computed during the initial training phase (see Line 11 of Algorithm 2) and use it as a stopping criterion. To address this inconvenience, the number of iterations to perform is provided to mP2L as a message $\mu_{2}=(\mu_{2}^{1},\ldots,\mu_{2}^{T})$. That is, for each task $t$, with a buffer $\mathcal{B}$, the message component $\mu_{2}^{t}$ is chosen in $\{1,\ldots,n_{t}+|\mathcal{B}|\}$, giving the number of mP2L’s iterations to perform.
$\mathscr{M}(\operatorname{\mathbf{j}},T)\,=\,\{2,\ldots,T\}^{|\operatorname{\mathbf{j}}|}\times\left[\textstyle\bigtimes_{t=1}^{T}\{1,\ldots,n_{t}+|{\mathcal{B}}|\}\right].$ (7)
A.2 Reconstruction Function
Under the assumption that the input parameters $\theta_{0},\gamma,m,\omega$ are the same for the compression function (Algorithm 4) and for CoP2L (Algorithm 3), the composition of the compression and the reconstruction function must output the exact same predictor as CoP2L. However, it only has access to the compression sets and the messages. The first challenge of the compression function is addressed using $\mu_{1}$ from line 6 to 11 of Algorithm 5. For each datasets, we verify if a datapoint was previously excluded from the buffer by the sampling function. If so, we remove it from the buffer. The second challenge is addressed in line 4 of Algorithm 5 by giving the message $\mu_{2}$ to mP2L as stopping criterion.
Algorithm 5 Reconstruction function of Continual Pick-To-Learn (CoP2L)
0:  $S^{\operatorname{\mathbf{i}}}_{1},S^{\operatorname{\mathbf{i}}}_{2},\ldots,S^{\operatorname{\mathbf{i}}}_{T}$Compression sets}
0:  $S^{\operatorname{\mathbf{j}}}_{1},S^{\operatorname{\mathbf{j}}}_{2},\ldots,S^{\operatorname{\mathbf{j}}}_{T}$Compression sets}
0:  $\mu_{1,1},\mu_{1,2},\ldots,\mu_{1,T}${Message sets}
0:  $\mu_{2}${Message set}
@@ A.2 Reconstruction Function
3:    $\hat{S_{t}}\leftarrow\{({\mathbf{x}},y,1)\}_{({\mathbf{x}},y)\in S^{\operatorname{\mathbf{i}}}_{t}}$
@@
5:    $B_{t}\leftarrow(S^{\operatorname{\mathbf{i}}}_{t}\setminus C^{\star})\cup S^{\operatorname{\mathbf{j}}}_{t}$
@@
7:       for $(\boldsymbol{x}_{i,j},y_{i,j},\omega_{i})\in S^{\operatorname{\mathbf{j}}}_{i}$ do
@@
A.3 Computation of the Bound
To compute the bound of Theorem 3.1, we need three things : the size of the first compression set $\operatorname{\mathbf{i}}$, the size of the second compression set $\operatorname{\mathbf{j}}$ and the message $\mu$. By simply running our CoP2L, we are able to retrieve $|\operatorname{\mathbf{i}}|+|\operatorname{\mathbf{j}}|$, but not the size of each compression set individually. Moreover, we are completely unable to retrieve the messages. Thus, instead of implementing Algorithm 3, we actually implemented the compression function as described in Algorithm 4.
@@ A.3 Computation of the Bound
Moreover, we use a block version of P2L, as defined by Algorithm 2 of [23]. We thus add $k$ datapoints at once to the compression set at Line 6 of Algorithm 2 (the used values of $k$ for each experiments is given in Appendix G. Thus, the value of $\mu_{2}^{t}$ (Line 6 of Algorithm 4) becomes the number of iterations of the mP2L algorithm at for task $t$. Knowing that mP2L outputs a compression set $C^{\star}_{t}$, and adds $k$ datapoints to the compression set at each iteration, the number of iterations is $\mu_{2}^{t}=\frac{|C^{\star}_{t}|}{k}$. Finally, when computing the bound, we use the code provided by [65] to invert the $\mathrm{kl}$ divergence.
Appendix B Tighter sample compression $\mathrm{kl}$ bound
@@
In this section, we prove a tighter $\mathrm{kl}$ bound that is not derived from Theorem 2.1 but proved in a very similar way. To do so, we start by providing the Chernoff test-set bound [66] for losses in $[0,1]$ [67].
Theorem B.1 ([67]).
@@
$p\leq\mathrm{kl}^{-1}\quantity(\tfrac{1}{n}\textstyle\sum_{i=1}^{n}X_{i},\tfrac{1}{n}\log\tfrac{1}{\delta}).$
@@
We prove the complement of the expression in Theorem B.2. Denote the upper bound
@@
$U_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}(\delta)=\mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S}^{\,\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\quantity(\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu)),\frac{1}{n-\operatorname{|\mathbf{i}|}-|\operatorname{\mathbf{j}}|}\log\frac{1}{P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{i}})P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})P_{\mathscr{M}(\operatorname{\mathbf{j}})}(\mu)\delta})$
@@
$\operatorname*{\mathbb{P}}_{S\sim\operatorname{\mathcal{D}}^{n}}\quantity(\exists\operatorname{\mathbf{i}},\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n),\mu\in\mathscr{M}(\operatorname{\mathbf{j}}):\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu))>U_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}(\delta))$